# Navigating Data Corruption in Machine Learning: Balancing Quality, Quantity, and Imputation Strategies

### v1

@@ 3 Learning with corrupted data
Figure 4 illustrates the relationship between the data missing ratio and model performance. Both experiments (NLP-SL and Signal-RL) show a gradual initial decline in performance, followed by a steeper descent, culminating in a sharp drop-off as the data missing ratio approaches 1.0. For reference, we include the performance of an optimized fixed-timing signal as a benchmark in the figure. The RL-trained signal outperforms the fixed-timing signal when the data missing ratio is less than 0.8.
@@
Tables 1 and 2 provide the detailed data for these experiments. In Figure 5, we change the x-axis to represent 1 - corruption ratio and fit the curve to the function in Equation 1. The fitted parameters for the NLP-SL experiment are: $a=0.475$, $b=3.517$, where $b$ represents the decay rate that controls the curve’s steepness. The goodness-of-fit analysis results in $R^{2}=0.995$, indicating that the exponential cumulative distribution function (CDF) is an excellent model for the observed performance. Similarly, for the Signal-RL experiment, the fitted curve parameters are: $a=395.827$, $b=7.493$ with $R^{2}=0.956$.
@@
This striking coincidence reveals an important and universal rule in machine learning: the diminishing return of data. The decay rate $b$ reflects the task’s nature, with the RL task demonstrating a much larger decay rate. This implies that RL tasks are more sensitive to noise compared to NLP tasks.
$S=a(1-e^{-b(1-p)})$ (2)
@@
where parameter $a=\frac{S_{0}}{1-e^{-b}}$; and $S_{0}$ is the model score when $p$ = 0 (i.e., no corruption).
Here, we provide an explanation for the observed pattern. For both NLP and DRL tasks, the model relies on recognizing patterns (e.g., semantic patterns, queuing patterns) in the data. When the data is completely corrupted, the critical patterns necessary for performance are entirely lost. As corruption decreases, the model rapidly recovers key patterns, resulting in a steep improvement in performance. However, once most of the critical patterns have been recovered, the marginal utility of additional clean data diminishes, leading to a saturation of performance.
As we know, the probabilities of observing rare events over a large number of trials converge to the Poisson distribution. When $n$ (the number of trials) is large, $p{\prime}$ (the probability of success per trial) is small, and the expected number of successes $\lambda=n\cdot p{\prime}$ is finite, the binomial distribution approximates the Poisson distribution:
$P(K=k)\approx\frac{\lambda^{k}e^{-\lambda}}{k!},$ (3)
where $\lambda=n\cdot p{\prime}$ is the rate parameter, representing the expected number of successes.
he exponential term $e^{-\lambda}$ in the Poisson distribution describes the probability of observing 0 successes, or equivalently, the probability of failure. In both experiments, the dataset size $n$ is large, making the recovery of individual patterns a rare event that can be modeled using the Poisson distribution. In our experiments, each pattern has an equal probability $p$ of being corrupted, or a probability $x=1-p$ of being recovered. A pattern can be recovered multiple times in the dataset. The probability of failing to recover such a pattern in a dataset with corruption level $p$ is proportional to $e^{-x}$. This leads to a system where the rate of change in performance depends on the difference between the system’s current state and its limit. Such behavior can be modeled using Equation 4, whose solution corresponds to Equation 1:
$\frac{dS}{dx}=abe^{-bx}=b(a-S)$ (4)
where $x=1-p$.
This model effectively explains the exponential improvement observed in performance as corruption decreases and clean data proportion increases.
@@
The Signal-RL model scores under inserting-noise and masking-region types of data corruption are shown in Figure 6. From Figure 6(a), we observe that inserting noise is significantly more detrimental than data-missing. The model’s performance deteriorates much more rapidly as the noise level increases, falling below the fixed-timing signal performance as soon as the noise level exceeds 10%. Additionally, the model scores become unstable, as indicated by the oscillations in Figure 6(a). The training process also becomes unstable, as shown in Figure 7.
@@
Masking-region is a unique type of data-missing corruption. Information about vehicles farther away from the intersection is less critical. By gradually discarding less important data, we observe that the model score only experiences a sharp decline when the masking ratio p exceeds 0.7. This observation leads to an empirical rule: 30% of the data is critical and determines the model’s performance, while the remaining 70% can be discarded without significantly affecting the model’s performance.
Figure 1: Isolated intersection simulation environment (SUMO)
Figure 2: Model convergence of two experiments. (a) NLP-SL experiment (b) Signal-RL experiment
Figure 3: NLP-SL model performances under different data corruption ratios. Most tasks exhibit Equation 1 pattern except MRPC and RTE tasks.
Figure 4: Model performances under different data corruption ratios (a) NLP-SL experiment. (b) Signal-RL experiment.
Figure 5: The relationship between model learning and data corruption ratio fitted by Equation 1; X-axis is 1 - data corruption ratio. (a) NLP-SL experiment. (b) Signal-RL experiment. Both experiments exhibit diminishing returns when data quality improves.
Figure 6: Signal-RL model performance under varying data corruption ratios (a) corruption mode = inserting noise. Inserting noise is more detrimental than data missing; and model performance becomes unstable. (b) corruption mode = masking region. Model performance does not degrade when 70% less important data are lost.
Figure 7: Training process becomes unstable when corruption mode is inserting noise
@@
Table 1: NLP-SL model scores at different data missing ratios.
@@
Table 2: Mean and standard deviation of Signal-RL model scores at different data missing ratios.
@@ 4 Effectiveness of Data Imputation
In this section, we assess whether and how different imputation strategies mitigate the impact of missing data. We note that the decision on whether to impute missing data involves a trade-off between data-missing and noise-inserting. We distinguish two common types of data-missing scenarios often encountered in practice:
  1. 1.
    The first type occurs when the exact location of missing data is known. This is typical for NLP tasks, where missing words are marked with tokens like “[UNK].”
  2. 2.
    The second type occurs when the location of missing data is not known. For example, in the Signal-RL experiment, vehicles are sometimes not detected, and we do not know which elements of the state vector are corrupted. To impute data in this second scenario, every possible location must be checked, which introduces significantly more noise than in the first case.
We will refer to these two scenarios as “exact imputation” and “general imputation,” respectively.
Experiment Design
We experiment with these imputation methods under varying data-missing ratios. There are various types of imputation methods, as reviewed in the literature. However, evaluating the effectiveness of these algorithms is non-trivial, and their accuracies are not adjustable parameters. For our study, where we focus on the trade-off between data-missing and noise, it is crucial to control the accuracy of imputation. To achieve this, we propose using an “inserting-noise” imputation method. In this method, accurate words (for NLP-SL) or state elements (for Signal-RL) are randomly replaced with random values, allowing us to control the imputation noise level, denoted as $q$. Later, we also experiment with traditional imputation methods and evaluate their effectiveness.
Let S(p) represent the model score when the data-missing ratio is $p$ and no imputation is applied. Let $\tilde{S}(p,q)$ represent the model score when the data-missing ratio is $p$ and an imputation method with noise level $q$ is used to preprocess the data before training. We define the imputation advantage, $A(p,q)$, as:
@@
$A(p,q)$ quantifies the improvement (or harm) caused by imputation relative to no imputation. Figure 8 shows the heatmap of imputation advantage for both experiments.
Observations
Several interesting observations can be made from the heatmaps. First, the Signal-RL task is much more sensitive to imputation noise, primarily because it involves general imputation. Red regions on the heatmap indicate areas where imputation is beneficial, while blue regions indicate areas where imputation harms model performance. The heatmaps clearly demonstrate the trade-off between data-missing and noise, and we identify two distinct regions:
  1. 1.
    Red Corner: This region is located in the lower right corner of the heatmap, where the data-missing ratio is high, and the imputation noise level is low. Accurate imputation in this region restores critical information, leading to significant improvements in model performance.
  2. 2.
    Blue Edge: This region is near the edge where $q$ = 1. When the imputation noise level approaches 1.0, the noise introduced during imputation overwhelms the model, leading to performance degradation. Interestingly, the greatest harm occurs when the data-missing ratio is around $p$ = 0.6.
Additionally, we draw the contour line corresponding to $A(p,q)=0$ on the heatmap. This line represents the decision boundary between the imputation-advantageous and imputation-disadvantageous regions. A few notable differences between the Signal-RL and NLP-SL tasks can be observed:
  1. 1.
    Signal-RL: The contour curve for the Signal-RL task lies much lower and is shifted to the right compared to the NLP-SL task. Moreover, the contour curve for Signal-RL is more ragged and fits well to an exponential curve that starts at (0, 0) and intersects the line $p=1$.
  2. 2.
    NLP-SL: The contour line for the NLP-SL task is smoother and fits well to a logistic function. When $p$ in range $[0,0.4]$, the decision boundary is relatively stable and remains around $q=0.68$. For $p$ in range $[0.4,1.0]$, the contour line transitions into a sigmoid curve, with its midpoint around $p=0.7$. This gradual transition reflects the trade-off between recovering critical information through accurate imputation and introducing ambiguities (e.g., incorrect word predictions) through noisy imputation.
The sigmoid shape of the NLP-SL contour line reflects a smoother transition between advantageous and disadvantageous regions as imputation noise increases, whereas the Signal-RL task exhibits an exponential drop-off in performance due to the compounding effect of errors in sequential decision-making. We classify tasks based on their sensitivity to noise:
@@
Figure 8: Heatmap of imputation advantage; x-axis is corruption ratio $p$ and y-axis is imputation noise level $q$. (a) NLP-SL experiment; (b) Signal-RL experiment.
@@
Figure 9: Illustration of imputation advantage pattern. ”Red corner” and ”blue edge” are areas where imputation advantage and disadvantage concentrate respectively. Task with decision boundary below diagonal line is called ”noise-sensitive” and its decision boundary is exponential function shaped. Task with decision boundary above diagonal line is called ”noise-insensitive” and its decision boundary is Logistic function shaped.
@@
We also experimented with other common imputation methods. For NLP-SL, we designed two imputation methods: “wordvec” and “BERT.” The wordvec imputation method uses context word vector embeddings (e.g., GloVe) and cosine similarity to impute missing words. The BERT method leverages the pre-trained bert-base-uncased model for imputation. The results of these experiments are shown in Figure 11 (a) and (b). To speed up preprocessing, a subset ratio of subset_ratio = 0.1 was used for BERT imputation.
The results indicate that BERT imputation improves the model score, whereas the wordvec method actually harms performance, suggesting that the wordvec approach introduces more noise than it resolves. An important observation is that common imputation methods do not maintain a fixed accuracy level as the missing data ratio $p$ varies. Instead, their performance corresponds to curves, rather than horizontal lines, on the advantage heatmap.
For Signal-RL tasks, we designed an imputation method called “context-filling.” In this method, road occupancy cells with a value of zero are imputed as having a value of one if there are enough vehicles surrounding them, as illustrated in Figure 10. The model scores with context-filling imputation compared to no-imputation are shown in Figure 11. The results demonstrate that context-filling imputation introduces more harm than benefits for this noise-sensitive task.
The findings in this section provide valuable insights for making practical decisions about data imputation strategies. Specifically, they highlight the importance of tailoring imputation methods to the nature of the task and the type of data corruption present.
Figure 10: Illustration of context-filling imputation for Signal-RL experiment.
Figure 11: Effectiveness of other imputation methods.

### v2

@@ 3.2 Observations
Figure 3 illustrates the relationship between the data missing ratio and model performance. Both experiments (NLP-SL and Signal-RL) exhibit an initial gradual performance decline, followed by a steeper descent that culminates in a sharp performance drop as the data missing ratio approaches 1.0. For reference, the figure includes the performance of an optimized fixed-timing signal as a benchmark. The RL-trained signal outperforms the fixed-timing signal when the data missing ratio is below 0.8.
@@
Tables 2 and 3 provide the detailed data for these experiments. In Figure 4, the $x$-axis is changed to represent (1 - corruption ratio) and fit the curve to the function in Equation 2. The fitted parameters for the NLP-SL experiment are: $a=0.475$, $\lambda=3.517$, where $\lambda$ represents the decay rate that controls the curve’s steepness. The goodness-of-fit analysis results in $R^{2}=0.995$, indicating that the exponential cumulative distribution function (CDF) is an excellent model for the observed performance. Similarly, for the Signal-RL experiment, the fitted curve parameters are: $a=395.8$, $\lambda=7.493$ with $R^{2}=0.956$.
@@
This striking coincidence reveals an important and universal rule in machine learning: the diminishing return of data. The decay rate $b$ reflects the task’s nature, with the RL task demonstrating a much larger decay rate. This implies that RL tasks are more sensitive to data corruption compared to NLP tasks. An explanation for Equation 2 is provided at the end of this section.
$S=a(1-e^{-\lambda(1-p)})$ (2)
@@
where parameter $a=\frac{S_{0}}{1-e^{-\lambda}}$; and $S_{0}$ is the model score when corruption ratio $p$ = 0 (i.e., no corruption).
@@
The Signal-RL model scores under inserting-noise and masking-region types of data corruption are shown in Figure 5. From Figure 5(a), results show that inserting noise is significantly more detrimental than data-missing. The model’s performance deteriorates much more rapidly as the noise level increases, falling below the fixed-timing signal performance as soon as the noise level exceeds 10%. Additionally, the model scores become unstable, as indicated by the oscillations in Figure 5(a). The training process also becomes unstable, as shown in Figure 6.
@@
Masking-region is a unique type of data-missing corruption. Information about vehicles farther away from the intersection is less critical. By gradually discarding less important data, it’s observed that the model score only experiences a sharp decline when the masking ratio $p$ exceeds 0.7. This observation leads to an empirical rule: 30% of the data is critical and determines the model’s performance, while the remaining 70% can be missing without significantly affecting the model’s performance. The exact numbers may not hold for other tasks. And for many tasks (e.g. previous NLP tasks) this kind of screening is not feasible.
Figure 3: Model performance across varying data missing ratios. The $x$-axis indicates the data missing ratio, while the $y$-axis shows the model score—normalized accuracy for NLP-SL and normalized negative queue length for Signal-RL. (a) NLP-SL experiment; (b) Signal-RL experiment. In both cases, model performance initially declines gradually, followed by a steeper drop and a sharp collapse as the missing ratio approaches 1.0. The RL-trained signal outperforms the fixed-timing signal when the data missing ratio is less than 0.8.
[…]
Figure 5: Signal-RL model performance under varying data corruption ratios. (a) Noise insertion: The $x$-axis represents the noise level. Inserting noise is more detrimental than missing data, often leading to unstable model performance. (b) Masked region: The $x$-axis denotes the masking-region ratio. A ratio of $p$ indicates that the farthest $400\times p$ meters of each lane are hidden from the model, simulating the limited visual range of video-based vehicle detectors. Model performance remains stable even when up to 70% of less critical data are removed.
Figure 6: Training instability under noise-insertion corruption. The $x$-axis represents training episodes, and the $y$-axis shows episode return. From (a) to (d), the observation noise ratio increases from 0.0 to 0.4, leading to progressively more unstable training dynamics.
@@
Table 2: NLP-SL model scores at different data missing ratios; plotted in Figure 3(a).
@@
Table 3: Mean and standard deviation of Signal-RL model scores at different data missing ratios; plotted in Figure 3(b).
3.3 Explanation
An explanation for the observed pattern in Figure 4 and Equation 2 is provided here. For both NLP and DRL tasks, the model relies on recognizing patterns (e.g., semantic patterns, queuing patterns) in the data. When the data is completely corrupted, the critical patterns necessary for performance are entirely lost. As corruption decreases, the model rapidly recovers key patterns, resulting in a steep improvement in performance. However, the marginal utility of additional clean data diminishes, leading to a saturation of performance.
First, some basic probability theory is introduced. The probabilities of observing rare events in a large number of trials converge to the Poisson distribution [6]. When $n$ (the number of trials) is large, $p^{\prime}$ (the probability of success per trial) is small and the expected number of successes $\lambda=n\cdot p^{\prime}$ is finite, the binomial distribution approximates the Poisson distribution:
$P(K=k)\approx\frac{\lambda^{k}e^{-\lambda}}{k!},$ (3)
where $\lambda=n\cdot p^{\prime}$ is the rate parameter, representing the expected number of successes. The exponential term $e^{-\lambda}$ in the Poisson distribution describes the probability of observing 0 successes.
Next, the above theory is applied to our pattern recognition problem. In both experiments, the dataset size $n$ is large, making the discovery of a pattern from an individual sample a rare event. In our experiments, each pattern has an equal probability $p$ of being corrupted. Let $x=1-p$. A pattern can be recovered multiple times from samples in the dataset with probability $P(K=k)$. The probability of failing to recover such a pattern with corruption level $p$ is $e^{-\lambda x}$, where $\lambda$ is the pattern appearance rate given no corruption. As $x$ increases, the probability of failing to recover a pattern have derivative $-\lambda e^{-\lambda x}$. In other words, the probability of identifying such a pattern increase by $\lambda e^{-\lambda x}$ marginally. Suppose that model socre $S$ is proportional to the number of patterns identified. Then, as $x$ increases, the rate of $S$ is proportional to $\lambda e^{-\lambda x}$. This leads to Eqn 4 where $a$ is coefficient for this linear relation. Its solution corresponds to Eqn 2. It can be shown that $\lambda e^{-\lambda x}=\lambda(a-S)$, so Eqn 4 actually describes a dynamic system where the rate of change in performance depends on the difference between the system’s current performance and its limit.
$\frac{dS}{dx}=a\lambda e^{-\lambda x}$ (4)
where $S$ is the model score; $x=1-p$ and $p$ is corruption rate; $b$ is pattern appearance rate; $a$ is the model score when there is no corruption.
@@ 4 Effectiveness of Data Imputation
This section is aimed at assessing whether and how different imputation strategies mitigate the impact of missing data. It is noted that the decision to impute missing data involves a trade-off between recovering missing information and potentially introducing noise.
4.1 Experiment Design
These imputation methods will be evaluated under varying data-missing ratios. There are various types of imputation methods, as reviewed in the literature. However, evaluating the effectiveness of these algorithms is non-trivial, and their accuracies are not adjustable parameters. For this study, where the focus is on the trade-off between data-missing and noise, it is crucial to control the accuracy of imputation. To achieve this, an artificial imputation method - “inserting-noise” - is proposed. For this method, accurate words (for NLP-SL) or state elements (for Signal-RL) are randomly (with probability $q$) replaced with random values, allowing us to control the imputation method noise level $q$. The example code for NLP-SL is provided below. The function perturb_sentence is responsible for adding missing-type corruption to clean data and carrying out the artificial imputation. Later in this section, traditional imputation methods are also evaluated.
Let $S(p)$ represent the model score when the data-missing ratio is $p$ and no imputation is applied. Let $\tilde{S}(p,q)$ represent the model score when the data-missing ratio is $p$ and an imputation method with noise level $q$ is used to preprocess the data before training. The imputation advantage, $A(p,q)$, is defined as:
@@ 4.1 Experiment Design
$A(p,q)$ quantifies the improvement (or harm if negative) caused by imputation relative to no-imputation. Figure 7 shows the heatmap of imputation advantage for both experiments.
4.2 Observations
Several interesting observations emerge from the heatmaps. First, the Signal-RL task demonstrates significantly greater sensitivity to imputation noise. This can be explained by the fact that sequential decision-making is inherently more noise-sensitive, combined with this task’s use of “general imputation”. In the heatmaps, red regions indicate where imputation improves model performance, while blue regions show where imputation is detrimental. These visualizations clearly reveal the trade-off between data missingness and noise introduction, with two distinct regions being identifiable:
  • Imputation advantageous corner: This region is located in the lower right corner of the heatmap, where the data-missing ratio is high, and the imputation noise level is low. Accurate imputation in this region restores critical information, leading to significant improvements in model performance.
  • Imputation disadvantageous edge: This region is near the edge where $q$ = 1. When the imputation noise level approaches 1.0, the noise introduced during imputation overwhelms the model, leading to performance degradation. Interestingly, the greatest harm occurs when the data-missing ratio is around $p$ = 0.6.
Additionally, the black dashed contour line corresponding to $A(p,q)=0$ is overlaid on the heatmap, indicating the decision boundary between regions where imputation is advantageous versus disadvantageous. The black solid line shows the fitted decision boundary. The green and lime green dashed lines denote the 68% and 95% confidence intervals, corresponding to ±$\pm 1$ and ±$\pm 1.96$ standard errors, respectively. A few notable differences between the Signal-RL and NLP-SL tasks can be observed:
  • Signal-RL Decision Boundary: The contour curve for the Signal-RL task lies much lower and is shifted to the right compared to the NLP-SL task. Moreover, the contour curve for Signal-RL is more ragged and fits to an exponential curve that starts at (0, 0) and intersects the line $p=1$.
  • NLP-SL Decision Boundary: The contour line for the NLP-SL task is smoother and fits well to a logistic function. When $p$ in range $[0,0.4]$, the decision boundary is relatively stable and remains around $q=0.68$. For $p$ in range $[0.4,1.0]$, the contour line transitions into a sigmoid curve, with its midpoint around $p=0.7$. This gradual transition reflects the trade-off between recovering critical information through accurate imputation and introducing ambiguities (e.g., incorrect word predictions) through noisy imputation.
The sigmoid shape of the NLP-SL contour line reflects a smoother transition between advantageous and disadvantageous regions as imputation noise increases, whereas the Signal-RL task exhibits an exponential drop-off in performance due to the compounding effect of errors in sequential decision-making. tasks can be classified based on their sensitivity to noise:
@@ 4.2 Observations
Figure 7: Heatmap of imputation advantage. The $x$-axis denotes the data missing ratio $p$, and the $y$-axis represents the imputation noise level $q$. The black dashed line indicates the decision boundary separating regions where imputation is beneficial from those where it is detrimental. Black solid line is the fitted curve of decision boundary. The green and lime green dashed lines denote the 68% and 95% confidence intervals. (a) NLP-SL experiment: NLP-SL is a noise-insensitive task. The decision boundary lies above the diagonal and is well approximated by a logistic function. (b) Signal-RL experiment: Signal-RL exemplifies a noise-sensitive task. The contour line appears below the diagonal, more irregular in shape, and fits an exponential curve starting from (0, 0) and intersecting the line $p=1$.
@@
Figure 8: Illustration of imputation advantage pattern. The $x$-axis denotes the data missing ratio $p$, and the $y$-axis represents the imputation noise level $q$. “Imputation advantageous corner” and “Imputation disadvantageous edge” are areas where imputation advantage and disadvantage concentrate respectively. Task with decision boundary below diagonal line is called “noise-sensitive” and its decision boundary is fitted by exponential function. Task with decision boundary above diagonal line is called “noise-insensitive” and its decision boundary is fitted by Logistic function.
@@
Additional experiments were conducted using other common imputation methods. For NLP-SL, two imputation methods are introduced: “wordvec” and “BERT”. The wordvec imputation uses context word vector embeddings (GloVe) and cosine similarity to impute missing words. The BERT method leverages the pre-trained bert-base-uncased model for imputation. These two imputation methods are examples of “exact imputation”. The results of these experiments are shown in Figure 10 (a) and (b). To speed up preprocessing, a subset ratio of 0.1 was used for BERT imputation.
The model score for the NLP-SL task shows an increasing standard error, rising from 0.0015 to 0.0227 as the data corruption ratio increases from 0.1 to 1.0. Imputation using Word2Vec leads to a significant performance decline relative to the no-imputation baseline (Figure 10(a)), indicating that this approach introduces more noise than it mitigates. BERT-based imputation effectively reconstructs informative content for classification while introducing minimal additional noise, resulting in better performance compared to the no-imputation condition (Figure 10(b)). At a noise-missing ratio of 0.3, the model score difference between BERT-based imputation and the no-imputation baseline is +0.046, with a standard error of 0.0043, yielding a $z$-score of 10.65. This difference is statistically significant. An important observation is that common imputation methods do not maintain a fixed accuracy level as the missing data ratio $p$ varies. Instead, their performance corresponds to curves, rather than horizontal lines, on the advantage heatmap.
For Signal-RL tasks, an imputation method called “context-filling” is proposed. In this approach, road occupancy cells with a value of zero are imputed as one if sufficient surrounding vehicles are detected, as illustrated in Figure 9. This imputation method is an example of “general imputation”, where each element of the state vector is evaluated for potential correction. Model performance with context-filling imputation, compared to the no-imputation baseline, is presented in Figure 10(c). The model score exhibits substantial variability, with standard errors ranging from 4.8 to 47.6 as the corruption level increases from 0.1 to 0.9. Context-filling shows no clear advantage and in fact performs slightly worse in our results. At a data-missing ratio of 0.6, the model score between context-filling and no imputation is 22, with a standard error of 15. The $z$-score is 1.48 and the $p$-value is 0.14. This difference is not substantial given this high performance variance observed in the Signal-RL experiments.
Figure 9: Illustration of context-filling imputation in the Signal-RL experiment. This represents a type of “general imputation”, where each element of the state vector is evaluated for potential correction. Road occupancy cells with a value of zero are imputed as one if sufficient surrounding vehicles are detected.
Figure 10: Effectiveness of alternative imputation methods. The $x$-axis represents the data missing ratio, and the $y$-axis indicates the model score (normalized accuracy). Solid lines denote the no-imputation baseline, while dashed lines represent model performance with imputation. (a) Imputation using word vectors (NLP-SL task): This method introduces substantial noise, resulting in a significant performance drop compared to the no-imputation baseline. (b) Imputation using BERT (NLP-SL task): BERT-based imputation effectively recovers informative content useful for classification while introducing considerably less noise, leading to improved performance over the no-imputation baseline. At $p$ = 0.3, model score difference = +0.046, std error = 0.0043, $z$-score = 10.65; the difference is significant. (c) Context-filling imputation (Signal-RL task): This method shows no clear advantage over the no-imputation. At $p$ = 0.6, model score difference = -22, std error = 15, $z$-score = -1.48, $p$-value = 0.14; the difference is not significant.